OpenStudy (anonymous):

-2sin^2(x)-1cos(x)=-1

OpenStudy (anonymous):

The smallest non-negative solution is: The next smallest non-negative solution is:

OpenStudy (campbell_st):

sin^2(x) = 1 - cos^2(x) use the substitution -2(1 - cos^2(x)) - cos(x) = -1 2cos^2(x) - cos(x) - 1= 0 (2cos(x) +1)(cos(x) -1)=0 cos(x) = -1/2 and 1 next you need to find the angles that represent cos(x) = -1/2 and cos(x) = 1 it will depend on the domain if its 0<= x<= 360 then there will be 3 answers

OpenStudy (anonymous):

What's it talking about when it asks me what the smallest non-negative solution and the next smallest non-negative solution? It doesn't provide a domain/range.

OpenStudy (anonymous):

i wanna to ask a question is there anyone here heard abt. the double angle!

OpenStudy (campbell_st):

the non negative solutions occur when x>= 0 then the solution is x = 0, 360, 720 etc...